if any wave is reflected by an obstacle and ratio of amplitude of antinode to the node is 8. the percentage of energy that is transmitted is?

Grade:12

1 Answers

praneeth bonu
12 years ago

A1+A2/A1-A2=8

A1/A2=9/7

ENERGY PROP. TO AMPLITUDE^2

SO,% of energy transmitted =[9/7]^2multiplyed by 100